OpenStudy (joshiscool):

Which value of y makes the equation true? y – 6.3 = –2.2 y =

OpenStudy (joshiscool):

I will medal and Fan

OpenStudy (youngstudier):

Are there answer choices or no?

OpenStudy (joshiscool):

no

OpenStudy (youngstudier):

Ok, so do you know what the first step to solving this equation would be?

OpenStudy (joshiscool):

http://prntscr.com/avm2is

OpenStudy (youngstudier):

@Joshiscool what is the first step we should take to solve this equation? :)

OpenStudy (anonymous):

The first, and in this case the only one, is to do reverse order of operations to isolate the variable So what would the first step be?

OpenStudy (joshiscool):

IDK

OpenStudy (youngstudier):

Ok, so first we want to isolate y. Do you know how to do that? :)

OpenStudy (anonymous):

Well if you have to do reverse order of operations then you would have to do the very last step of order of operations right? Which is addition/subtraction What could you add or subtract to isolate the variable?

OpenStudy (joshiscool):

Maybe the answer will tell me all by it's self and maybe I can give out medals to the one who tells me first

OpenStudy (anonymous):

....

OpenStudy (anonymous):

yeah no

OpenStudy (ashely14):

cant you just subtract -6.3 from -2.2

OpenStudy (youngstudier):

We're not allowed to give out direct answers, sorry.

OpenStudy (anonymous):

Yup Ashely14 got it

OpenStudy (ashely14):

i didnt give out a direct answer

OpenStudy (anonymous):

But remember not just from the -2.2 but you have to subtract from both sides Yeah dont worry thats not direct

OpenStudy (ashely14):

okay

OpenStudy (joshiscool):

I got 8.5

OpenStudy (ashely14):

so if you subtract them both you will get -4.1

OpenStudy (ashely14):

i think

alones (alones):

y−6.3=−2.2 add 6.3 to both sides. y−6.3+6.3=−2.2+6.3 y=4.1 Was thAT HELpful, Josh you need to understand The steps,

OpenStudy (joshiscool):

I added then I subtracted and I got the same number both times

alones (alones):

You made mistake in adding or maybe while you where subtracting

OpenStudy (ashely14):

I think another way he could do it is by (-6.3)-(-2.2)= -4.1